OpenStudy (anonymous):

x^2-49=0 by completing square

OpenStudy (anonymous):

Tried anything yet?

OpenStudy (kainui):

I don't think completing the square makes much sense here honestly lol.

OpenStudy (anonymous):

Well you can use it, x^2 = 49, x = 7, x = -7 not very helpful lol.

OpenStudy (kainui):

You can simply add 49 to both sides and then take the square root. Problem solved, no reason to over complicate this I think lol.

OpenStudy (anonymous):

(x^2-a^2) = (x+a)(x-a)

OpenStudy (anonymous):

Complete the square to solve for x. x2−49=0

OpenStudy (jdoe0001):

@mjacoby16 do you know what a "perfect square trinomial" is?

OpenStudy (anonymous):

Whenever you multiply a binomial by itself twice, the resulting trinomial is called a perfect square trinomial